Derrick is creating a model of the Eiffel Tower. The Eiffel Tower is 301 meter tall and has a base of width 100 meters. His model is will have a base of 10 cm. What will be the height of the model?
Answer by TimothyLamb(4379)

scale = m/r = 10/100 cm/meter
---
m/301 = 10/100
m = 301*10/100
m = 30.1 cm
---
model height = 30.1 cm
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---